vidyo

vidyo link

Action Items:

News:

  • Sew Yan will be interviewed for PhD position (for foreign students) in Padova this afternoon.
  • next meetings scheduled (see Franco's e-mail)
  • status report of ttH searches on CMS will be provide at the beginning of June (8th?)

Framework e produzione

  • spark cluster set up by Paul (config, etc..).
  • image to that created by Federica. Lots of problem to do it. No lustre up to now. Test today.
  • meeting tomorrow afternoon at 2:30pm to start discussing about future framework usage.

Attivita' di analisi

B physics

  • Paolo will meet with A.Bragagnolo to try to restart production for new analysis.

Dark Matter

Dibosoni, VZ risonante (V=W,Z)

  • analysis pre-approved. ARC formed.

hh4b non risonante

  • pre-approval talk done 2 weeks ago.
  • ARC formed (Rizzi is chair).
  • still waiting for official pre-approval since further checks were required about BDT validation strategy.
  • Pheno and LHCP missed. Target is EPS now.

Same sign leptons

Seesaw a 13 TeV

EPR

  • Pablo will work on BTV group.
  • Martino will get EPR from pixel commissioning work plus pixel DOC shifts.

Isolamento muoni

Flavor tagging

Trigger

Detector

A.O.B.:

-- MarcoZanetti - 2016-02-01

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2017-05-09 - MartinoDallOsso
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback